OpenStudy (anonymous):

A 10" diameter pumpkin pie is cut into six equal servings. What is the area of the top of each piece of pie?

OpenStudy (saifoo.khan):



OpenStudy (anonymous):

So to get to degree you times 10X6

OpenStudy (phi):

Have you memorized the formula for the area of a circle? I would use that. Then it seems each piece would be 1/6 of the total area.

OpenStudy (anonymous):

mmm.... pumpkin pie.... :)

OpenStudy (saifoo.khan):

i thought u like tarts more. -_-

OpenStudy (anonymous):

Geometry is hard plus if your home schooled

OpenStudy (phi):

Try to find the area of the circle.

OpenStudy (anonymous):

\[60/360*25 \]

OpenStudy (saifoo.khan):

you forgot the pi.

OpenStudy (anonymous):

ya and that equal \[4.16666667\]

OpenStudy (anonymous):

Thx for help

OpenStudy (phi):

not so fast.... A= pi r^2 you need to multiply by pi (about 3.14 or 22/7)

OpenStudy (anonymous):

I got these three choices 3.33inches 4.17inches 16.66inches

OpenStudy (phi):

Well, first it should say square inches or \(inches^2 \) for the units. The area for the entire pie is 3.14159*5*5= 78.54 sq in 1/6 of that (one piece) = 78.54/6= 13.09 sq in

OpenStudy (anonymous):

I have computer work and i choose 4.17 and it show me that the correct answer

OpenStudy (phi):

Out of curiosity, what program/school are you using?

OpenStudy (anonymous):

Oddesey

OpenStudy (phi):

Do you have access to a real person to show them this question? Unless the pumpkin pie is not round, that is not the correct answer, and you should ask somebody there to check this question. Otherwise you will have trouble learning this stuff....

OpenStudy (anonymous):

Well i can always ask the math teacher

OpenStudy (anonymous):

but in other way I wont need geometry later at least thats what i think

OpenStudy (anonymous):

but thx for your help

OpenStudy (phi):

First, if there is a problem it would be good for them to fix it (don't you think?) And learning this stuff is helpful, just the way learning how to dance or read is helpful. You might not notice, but you are learning how to think, and that is always a good thing to know how to do...

OpenStudy (anonymous):

That's true